In a study of monthly salary distribution of residents in Paris conducted in year 2015, it was found that the salaries had an average of €2200 (EURO) and a standard deviation of €550. Assume that the salaries were normally distributed.
Consider sampling with sample size 64 on the above population. Compute the mean of the
̅ sampling distribution of the mean (
1
Expert's answer
2020-04-06T17:57:16-0400

The sampling distribution of the mean from a normal distribution is normally distributed with mean "\mu" and standard deviation "\frac{\sigma}{\sqrt{n}}" . Thus mean of the sampling distribution of the mean is €2200 and standard deviation is "\frac{€550}{\sqrt{64}}=€68.75" . Thus the sampling distribution of the mean is N(€2200,€68.75).